OpenStudy (anonymous):

Susan is writing a linear equation for the cost of her cell phone plan. In the first month, she talks for 52 minutes and is charged $19.41. In the second month, she talks for 380 minutes and is charged $45.65. A) How much money does Susan's cell phone company charge for each minute? B) Write an equation that Susan can use to determine the cost of her cell phone plan, y, as it relates to the number of minutes used, x.

OpenStudy (anonymous):

what grade is this

OpenStudy (anonymous):

11th

OpenStudy (anonymous):

ok may be I know how to do this let me try this

OpenStudy (anonymous):

I find some answer to this

OpenStudy (anonymous):

52 19.41 380 45.65 --- slope = m = dy/dx m = (45.65 - 19.41)/(380 - 52) m = 0.08 --- y - 45.65 = 0.08(x - 380) y - 45.65 = 0.08x - 380*0.08 y = 0.08x - 380*0.08 + 45.65 y = 0.08x + 15.25 --- answer: A) charge for each minute = 8 cents B) y = 0.08x + 15.25

OpenStudy (anonymous):

I hope u pass

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

ur welcome